OpenStudy (anonymous):

1. what is the degree of the monomial 15x^3y^9? (1 point) 12 3 9 15

OpenStudy (anonymous):

soory about that

OpenStudy (cwrw238):

where you have 2 variables the degree is the highest sum of the powers in this case it is 3+9

OpenStudy (cwrw238):

the degree of 2xy^2 + 4xy would be 2+1 = 3

OpenStudy (anonymous):

2. what is the difference? 10c^3-8c^3 (1 point) 18c^3 2c^3 3c^6 -80c^6

OpenStudy (anonymous):

so the awnser for #1 was 3

OpenStudy (cwrw238):

no add the powers of x^3 and y^9 3 + 9 = 12

OpenStudy (anonymous):

its 12

OpenStudy (cwrw238):

as for you second question 10 eggs - 8 eggs = 2 eggs imagine c^3 = 1 egg!!!!

OpenStudy (anonymous):

so #1 is 12

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

then , what is number two

OpenStudy (cwrw238):

just look at my last post

OpenStudy (anonymous):

so its D for number 2 is -80*c^6

OpenStudy (anonymous):

which grade are you in lol just curious?

OpenStudy (cwrw238):

no look at the following subtractions: 2a - a = a 4b - 2b = 2b 6x - 2x = 4x 6c^3 - 2c^3 = 4c^3

OpenStudy (cwrw238):

i am a maths student in a british university

OpenStudy (cwrw238):

studying for a maths degree

OpenStudy (anonymous):

awsome you my biggest fan now and i truely mant that

OpenStudy (anonymous):

i ment to say "ment" lol

OpenStudy (cwrw238):

ty i gotta go now drews

OpenStudy (anonymous):

ok go on now you got work to do